Partial-year depreciation

Equipment acquired at a cost of $73,000 has an estimated residual value of $4,000 and an estimated useful life of 8 years. It was placed in service on May 1 of the current fiscal year, which ends on December 31. This information has been collected in the Microsoft Excel Online file. Open the spreadsheet, perform the required analysis, and input your answers in the questions below.

  1. Determine the depreciation for the current fiscal year and for the following fiscal year by the straight-line method. Round your answers to the nearest dollar.

    Depreciation
    Year 1 $  
    Year 2 $  
  2. Determine the depreciation for the current fiscal year and for the following fiscal year by the double-declining-balance method. Do not round the double-declining balance rate. Round your answers to the nearest dollar.

    Depreciation
    Year 1 $  
    Year 2 $  
Answer #1

a) straight line method

depreciation = cost - residula vale / useful life

= 73,000 - 4,000 / 8

= 8,625 per year

current fiscal year ( from 1st may to 31st december)

depreciation = 8,625 x 8/12 = 5,750

following fiscal year ( from 1st january to 31st december)

1st january to 30th april

8,625 x 4/12 = 2,875

1st may to 31st december

8,625 x 8/12 = 5,750

total depreciation = 8,625

depreciation
year 1 5,750
year 2 8,625

b) double-declining-balance method

depreciation rate = 100/useful life x 2

= 100/8 x 2

= 25%

depreciation for 1st year = 73,000 x 25% = 18,250

current fiscal year ( from 1st may to 31st december)

depreciation = 18,250 x 8/12 = 12,167

following fiscal year ( from 1st january to 31st december)

1st january to 30th april

18,250 x 4/12 = 6,083

depreciation for 2nd year = 54,750 x 25% = 13,688

1st may to 31st december

13,688 x 8/12 = 9,125

total depreciation = 15,208

depreciation
year 1

12,167

year 2 15,208